# Scattering Theory of Chiral Edge Modes in Topological Magnon Insulators

### v1

@@ Abstract
The scattering phase $\phi$ as a function of the relative momentum $q$ at different total momenta $K$ and anisotropies $\Delta$ are shown in Fig. 3 b) for interchain hopping $J_{y}=0.5$. As chiral magnons are strongly localized and thus predominately live on the outmost chain, it is pertinent to compare their scattering properties with those of magnons in a single one-dimensional XXZ chain [40]. While in some cases this is a good approximation, for certain values of $K$ and $\Delta$ we observe a strong renormalization. In particular, a comparison of the numerical results with analytical predictions for the one-dimensional XXZ chain at the edge (gray lines) reveals significant renormalization of the interacting theory at large momenta $K$ and small anisotropies $\Delta\in\{0.1,0.3\}$. In contrast, for small $K$ or large $\Delta$ the results align well with the bare XXZ predictions. It is worth noting that we find real-valued $\phi$ only indicating that there is indeed no weight of the wavefunction transferred to the bulk during scattering.
@@
To provide a more comprehensive comparison between the effective scattering theory and the bare XXZ model, we consider the momentum-dependent scattering length $a(k)\equiv\lim_{p\to k}\partial_{p}\phi(p,k)$ renormalized with the effective mass $m^{*}(k)\equiv(\partial_{k}^{2}\varepsilon(k))^{-1}$, and define the interaction $c(k)\equiv-1/[m^{*}(k)a(k)]$. The parameter $c(k)$ represents the contact interaction in an effective one-dimensional edge Hamiltonian for two magnons with small relative momenta: see Supplementary Material for details. In Fig. 3 c) we show the effective interaction as a function of the COM momentum for different anisotropies, observing strong deviations from the bare XXZ for large and positive momenta. As we will demonstrate, the observed strong renormalization effects in the edge theory can be traced back to bound modes in the bulk, which can resonate with the chiral magnon pair.
@@
Bulk Resonances – The energy of scattering edge states is determined by the one-magnon band and remains independent of $\Delta$ due to the large spatial separation between the two magnons of an asymptotic state. Bulk modes, however, can resonate with the scattering states. A sketch of the relevant nondispersing eigenstates, which describe two bulk magnons in close proximity and are thus strongly interacting, is shown in Fig. 3 a). Each chain contributes one strongly bound mode, which is highly sensitive to changes in the anisotropy $\Delta$. As interactions increase, these modes are eventually pushed out of the two-particle continuum, forming stable bound states (see Supplementary Material for more details) [49, 50]. The strong renormalization of scattering properties can be understood by considering an asymptotic state of two chiral magnons with total (and relative) momentum $K$ ($q$) within the continuum of scattering states, as illustrated in Fig. 4 a). As the anisotropy $\Delta$ increases, it lowers the energy of all bulk-bound modes, causing a series of resonances at interaction values $\Delta_{1}(K,q),...,\Delta_{N}(K,q)$. This effect strongly depends on the total momentum $K$ of the asymptotic states: For small $K$, all relevant bulk modes are already situated at lower energies than the asymptotic energy window, even at $\Delta=0$, as shown in Fig. 3 a). The situation is different though for large $K$.
@@
The number of resonant bulk modes $N$ is expected to remain finite also in the thermodynamic limit $L_{y}\to\infty$, as only bulk modes near the edges have significant overlap with edge modes. Once the highest-energy bulk mode moves out of resonance with the asymptotic continuum, edge scattering becomes well-described by the one-dimensional XXZ model, as seen in Fig. 3 b). We further illustrate this behavior in Fig. 4 b) - f), by examining how the characteristics of bulk-bound states change with varying $\Delta$. As $\Delta$ decreases from strong interactions, the energy of certain bulk-bound states enters the two-chiral magnon continuum, allowing them to hybridize with edge states. This hybridization is detected through the edge participation factor $P=\sum_{x}|\psi(x,y=1)|^{2}$, which shows three distinct peaks, signaling strong coupling with different asymptotic scattering states, as shown in Fig. 4 b) and c).
@@
We then compare the unhybridized wave function, shown in Fig. 4 d), with its form at two of the most prominent resonances, as depicted in Fig. 4 e) and f). In all cases, the wave function remains sharply peaked at small relative distances $x$, but at resonance, a plane wave profile emerges in the magnon distribution near the edge, indicating hybridization with scattering states. The resonances change the effective magnon interactions, which confirms the behavior seen in Fig. 3 c). Notably, when comparing the hybridized cases in Fig. 4 e) and f), additional nodes in the edge plane wave profile become apparent, which correspond to different relative momenta $q$ of the hybridizing asymptotic states. As the system size increases along the $x$-direction, the spectrum of asymptotic edge states becomes denser, characterized by a continuous range of relative momenta $q$. Consequently, discrete bulk modes are tuned through a continuum of edge modes. However, the extent of hybridization between bound bulk states and edge modes ultimately depends on the spatial overlap of their respective wavefunctions.
Discussion. – We have investigated the scattering of chiral edge magnons in a two-dimensional magnet subjected to an electric field gradient. By thoroughly analyzing the interactions among edge modes, we demonstrated their stability during scattering events and extracted their corresponding scattering phases. Much like Feshbach resonances in atomic physics, the scattering of chiral magnons is significantly renormalized through resonances with bulk modes. Our work opens the door to describing the many-body dynamics of edge modes within an effective one-dimensional framework, which in the two-magnons sector is governed by the effective Hamiltonian
where $U_{\text{2-eff}}(k_{1},k_{2})$ captures the effective interactions among the two chiral magnons of energies $\varepsilon(k_{1,2})$. The effective interaction $U_{\text{2-eff}}(k_{1},k_{2})$ is uniquely determined by the scattering phase $\phi(k_{1},k_{2})$ through methods from inverse scattering [51, 52, 53]. At coincident momenta, one recovers the effective interaction previously introduced $c(k)=U_{\text{2-eff}}(k,k)$, and Eq. (3) is mapped into the Lieb-Liniger model [54, 55] describing galilean bosons with contact interactions. A brief account for the effective Hamiltonian and its small momentum limit is reported in the Supplementary Material . It will be intriguing to use Eq. (3) as a building block for a full-fledged many-body effective description, and take advantage of both powerful analytical methods and advanced numerical techniques available in one dimension. Our study underscores the crucial influence of bulk spectra on the many-body behavior of edge states in topological magnon systems. It would be fascinating to observe the phenomenology we have presented in experiments, such as through the proposed injection-measurement protocol applied to topological magnon candidates such as CrI3 [29].
@@
Acknowledgements. – We acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC–2111–390814868, TRR 360 – 492547816, DFG grants No. KN1254/1-2, KN1254/2-1, and DFG Emmy Noether Programme—Project No. 504261060, the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 851161), as well as the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
@@
Data and Informations availability. – Data, data analysis, and simulation codes are available upon reasonable request on Zenodo [56].
Supplementary Information
Scattering Theory of Chiral Edge Modes in Topological Magnon Insulators
Stefan Birnkammer, Johannes Knolle, Michael Knap, Alexander Mook and Alvise Bastianello
@@
1 The Strong Anisotropy Limit – Chiral Magnon Bound States
@@
In the main text we emphasized how strongly bound eigenstates can be tuned through resonances with individual asymptotic chiral states. Having increased the bare interaction strength $\Delta$ to sufficiently large values these strongly localized modes will transition to form stable bound modes, separated in energy from the two-particle continuum, as shown in Fig. S1 a). In total a number of $L_{y}$ bound state bands will experience this effect resembling the results for the single magnon spectrum. Each bound state band, however, is associated with twice the shift in momentum space. This is a result of the AC phase coupling to the physical spin charge of the excitation, the latter is double for the bound state compared to individual magnon excitations. In the spirit of the single magnon spectrum sufficient coupling between the wires, nonetheless, opens a bulk gap stabilizing two chiral bound state modes at the edges of the system, see Fig. S1 b) - d). Same as the single magnon edge modes these edge bound states are stable throughout the tested parameter regime. This is consistent with findings from Ref. [49, 50].
2 Characterization of the Many-Body Theory of Edge Modes
1 Numerical Extraction of Scattering Phases
@@
Our investigations on the edge theory in the main text are primarily based on the momentum dependent analysis of scattering phases. For this, we rely on a protocol consisting of the following points to extract scattering information from the numerically obtained spectra, also summarized in Fig. S1 e).
@@
    Determine the momentum-dependent energy window $[E_{1}(K),E_{2}(K)]$ possible for pairs of chiral magnons from the single magnon spectrum.
@@
    Filter the eigenspectrum for the eigenenergies within $[E_{1}(K),E_{2}(K)]$.
@@
    From the fitted parameters $q,A,B$ compute the scattering matrix $S(q)=\dfrac{B}{A}$ as a function of relative momentum $q$.
@@
$S_{\mathrm{XXZ}}(K,q)=e^{i\phi_{\mathrm{XXZ}}(K,q)}\qquad\text{with}\qquad\phi_{\mathrm{XXZ}}(K,q)=\pi-i\log\Bigg{[}\dfrac{1-4\Delta e^{i(\frac{K}{2}+q)}+e^{iK}}{1-4\Delta e^{i(\frac{K}{2}-q)}+e^{iK}}\Bigg{]}.$ (S1)
A comparison between the XXZ predicted (bottom row) and the numerically extracted (top row) scattering phases $\phi$ is shown in Fig. S1 f) - h) for anisotropies $\Delta\in[0.0,0.4,1.0]$. We collect scattering phases for values of $K\in[-\dfrac{\pi}{4},\dfrac{\pi}{4}]$ and show the results as a function of the localization length $\ell$ of the eigenstates. To further discriminate the individual data points we highlight their relative momenta via color coding. Strongest deviations appear at small $\Delta$ for less localized states. This fits the semiclassical expectation of two more delocalized edge states having a reduced chance of scattering during their orbital motion at the boundary. In reality this enhanced leaking into the bulk, however, abets hybridization with bulk modes and the existence of resonances.
Figure S1: Chiral Bound States and Scattering Phase Analysis. a) Bound state bands separated in energy from the two magnon continuum for anisotropy $\Delta=1.25$ and uncoupled wires. b) - d) Increasing hopping strength between the wires $J_{y}\in\{0.25,0.50,0.75\}$ opens a bulk gap and stabilizes modes of chiral bound states. e) Analysis of the scattering phase consists of (1.) - (3.) selecting eigenstates within the chiral energy window for a given momentum sector $K$, (4.) fitting the states against a plane wave ansatz and (5.) compute the scattering matrix $S$. f) - h) Comparison of extracted (top row) and XXZ predicted (bottom row) scattering phases for different $K$ momenta and anisotropies $\Delta\in\{0.0,0.4,1.0\}$. The associated relative momenta are highlighted by color coding. Deviations are evident at low momenta ($\Delta=0.0$) and especially for less localized states.
In the main text we also extract the effective interaction for the scattering phase data. For this, we determine the slope of $\phi(q)$ in the linear regime around relative momenta $q=0$. We correct the results by additionally weighting them with the effective magnon mass $m^{*}$ obtained from the curvature of the single magnon dispersion, i.e. $m^{*}(K)=\dfrac{1}{\partial_{K}^{2}\varepsilon(K)}$. The result quantifies the one-dimensional scattering problem yielding the inverse product of effective magnon mass and the scattering length $a$ for the edge theory, which is described in the next section. To quantify deviations from the bare XXZ theory we extract the same quantity from the analytical XXZ scattering prediction (S1)
$c(K)=\Bigg{[}-\frac{1}{m^{*}a}\Bigg{]}(K)=\dfrac{4\Delta\cos\big{(}K\big{)}}{2\Delta-\cos\bigg{(}\dfrac{K}{2}\bigg{)}}.$ (S2)

### v2

@@ IV Characterization of Edge Scattering Phases
In practice, we focus on the two-magnon sector of the Hilbert space, consider the center-of-mass (COM) frame and numerically diagonalize the Hamiltonian for each total momentum $K$ with standard numerical methods, further detailed in the Appendix C.1. An example of the momentum-resolved low-energy continuum spectrum is shown in Fig. 3 (a) (gray shaded region) for parameters $(J_{y},\Delta)\!=\!(0.5,0)$. The topological bulk gap is visible in the momentum-integrated density of states (right panel). For our scattering analysis, we focus on the energy-momentum region allowed for two chiral magnons localized at the lower edge, as highlighted by the blue shaded area in Fig. 3 (a). The energy-momentum window $E_{1}(K),E_{2}(K)]$ for such an asymptotic state of two edge magnons with momenta $k_{1}$ and $k_{2}$ is determined from the chiral magnon dispersion, i.e. $E_{1}(K)\!\leq\!\varepsilon(k_{1})+\varepsilon(k_{2})\!\leq\!E_{2}(K)$, with $k_{1}\!+\!k_{2}\!=\!K$, where we find scattering states labeled by the relative momenta $q=(k_{1}-k_{2})/2$ and tabulate the scattering phase. The scattering phase $\phi$ as a function of the relative momentum $q$ at different total momenta $K$ and anisotropies $\Delta$ is shown in Fig. 3 (b) for interchain hopping $J_{y}=0.5$. As chiral magnons are strongly localized and thus predominately live on the outermost chain, it is pertinent to compare their scattering properties with those of magnons in a single one-dimensional XXZ chain (gray lines) [49]. Indeed, we find this to be a good approximation for most parameters, i.e. for small momenta $K\lesssim 0$ or large anisotropy $\Delta$. Interestingly, however, we find significant deviations of our numerical results from single chain analytical predictions for large momenta $K>0$ and small anisotropies $\Delta\in\{0.1,0.3\}$, indicating a strong renormalization of the interacting theory. It is worth noting that we find real-valued $\phi$ only, confirming that there is indeed no weight of the wavefunction transferred to the bulk during scattering.
@@
In Fig. 3 (c) we show the equal-momenta effective interaction $U_{\text{2-eff}}(K/2,K/2)$ as a function of the COM momentum $K$ for different anisotropies, observing strong deviations from the bare XXZ prediction for large and positive momenta. The observed strong renormalization effects in the edge theory can be traced back to bound modes in the bulk, which can resonate with the chiral magnon pair.
V Bulk Resonances
The energy of scattering edge states is determined by the one-magnon chiral band $\varepsilon(k)$. Depending on $\Delta$, different bulk modes can resonate with the scattering states. Two scattering magnons can form a metastable bound state: the longer the two magnons stay together, the larger their scattering shift will be. In principle, these metastable compounds could manifest for arbitrary velocities of the magnons, but we observe a strong signal only whenever their velocity is small. This suggests a hybridization with bulk bound states, which are not dispersive. It should be stressed that any bulk excitation has zero velocity in the thermodynamic limit, as semiclassically they move around cyclotron closed orbits with zero average velocity [54, 55].
@@ V Bulk Resonances
Near the perturbative regime of weakly-coupled chains, each chain contributes one such strongly bound mode. Crucially, although the number of bound state bands scales with $L_{y}$, the number of resonances $N$ stays finite in the thermodynamic limit $L_{y}\to\infty$ since only bound modes located close enough to the boundary can overlap with edge modes.
The bound state energy is highly sensitive to changes in the anisotropy $\Delta$. As interactions increase, these modes are eventually pushed out of the two-particle continuum, forming stable bound states (see Appendix A for details) [56, 57]. For small $K\lesssim 0$, all relevant bulk modes are already situated lower in energy than the asymptotic energy window, even at $\Delta\!=\!0$, see Fig. 3 (a), and so resonances are not possible. The situation is different though for large $K$.
@@
The emergence of resonances between edge and bulk bound modes for varying $\Delta$ in the regime of large $K$ is entailed in hybridization of the associated wavefunctions shown in Fig. 4 (b) - (f). As $\Delta$ decreases from strong interactions, the energy of certain bulk-bound modes enters the two-chiral magnon continuum, allowing them to hybridize with edge states, see Fig. 4 (a). This hybridization is detected through the edge participation factor $P\!=\!\sum_{x}|\psi(x,\!y\!=\!1)|^{2}$, which shows three distinct peaks at $\Delta\!\in\!\{0.245,0.32,0.36\}$, signaling strong coupling to different asymptotic scattering states, as shown in Fig. 4 (b) and (c).
@@
We then compare the unhybridized wave function, shown in Fig. 4 (d), with the hybridized wave function at two of the most prominent resonances, as depicted in Fig. 4 (e) and (f). In all cases, the wave function remains sharply peaked at small relative distances $x$, but at resonance, a plane wave profile emerges in the magnon distribution near the edge, indicating hybridization with scattering states. The resonances change the effective magnon interactions, which confirms the behavior seen in Fig. 3 (c). Notably, when comparing the hybridized cases in Fig. 4 (e) and (f), additional nodes in the edge plane wave profile become apparent, which correspond to different relative momenta $q$ of the hybridizing asymptotic states. As the system size increases along the $x$-direction, the spectrum of asymptotic edge states becomes denser, characterized by a continuous range of relative momenta $q$. Consequently, a discrete set of bulk modes is tuned through a continuum of edge modes. However, the extent of hybridization between bound bulk states and edge modes ultimately depends on the spatial overlap of their respective wave functions.
VI Experimental Relevance
Signatures of this interplay between bulk and edge modes can also be examined in experiment. Inelastic STS allows systematic studies of scattering between chiral magnons via a real-time injection measurement protocol, see Fig. 1 (a). Our protocol consists of 3 steps: (i) Inject in the sample two magnon excitations of given energy $E(k_{1}),E(k_{2})$ within the bulk-gap via spin-exchange interactions between tip and material [58, 36]. Both magnons are injected with a time delay $t_{0}$ directly relating to the distance between both excitations as $s_{0}=v(k_{1})t_{0}$, where $v(k_{1})$ denotes the velocity of the first excited slower magnon. (ii) After propagation both magnons will scatter after time $t^{*}=s_{0}/[v(k_{2})-v(k_{1})]$ and experience scattering shifts $d_{1}(k_{1},k_{2})$ and $d_{2}(k_{1},k_{2})$, as illustrated in Fig. 1 (b). (iii) After scattering both excitations are detected at time $t$ with distance $s(t)=[v(k_{2})-v(k_{1})](t-t^{*})+[d_{2}(k_{1},k_{2})-d_{1}(k_{1},k_{2})]$ using STM. Hence monitoring the positions of both chiral magnons after scattering experimentally using local measurements and comparing to the free evolution of individual magnons allows us to detect $d_{1},d_{2}$, see Fig. 1 (a). Insights from this setup tunable in the energy and velocity of the injected magnons can later be applied to more complex transport experiments. A more rigorous discussion of experimental caveats, including a comparison between the typically expected sizes of scattering shifts $d_{1},d_{2}$ and experimental limitations due to measurement resolution of current STM setups, is given in Appendix D. We emphasize that the most crucial requirement for the proposed experiment is the availability of a material realization of a topological magnon insulator with sufficiently long-lived edge states. While our construction scheme based on the AC effect enables a theoretical study of the relevant phenomena, experimental implementations are challenged by the weak coupling between spins and external electric fields. Consequently, strong electric field gradients are necessary, which may, in turn, induce additional intra-atomic effects. However, a more promising route was recently proposed in Ref. [40], where an array of STM tips is used to generate a skew-harmonic electromagnetic scalar potential, producing sizable electric-field gradients without the need for large overall electric-field amplitudes.
VII Discussion
We have investigated the scattering of chiral edge magnons in a two-dimensional magnet subjected to an electric field gradient. By thoroughly analyzing the interactions among edge modes, we demonstrated their stability during scattering events and extracted their corresponding scattering phases, which fully captures the effective interactions. It would be intriguing to use Eq. (1) as a building block for a full-fledged many-body effective description, and take advantage of both powerful analytical methods and advanced numerical techniques available in one dimension. For example, the long-wavelength dynamics of magnons with small relative momenta is that of one-dimensional bosons with contact interactions, i.e. the famous Lieb-Liniger integrable model [52, 53], whose thermodynamics [49] and hydrodynamics [59] can be analytically studied. Our study highlights the critical role of bulk spectra in shaping the many-body behavior of edge states in topological magnon systems. It would be exciting to see the phenomena we describe realized experimentally, for instance, through the proposed injection-measurement protocol. Promising platforms include topological magnon candidates like CrI3 [29], where recent STM experiments indicate signatures of in-gap edge magnons [60], or artificial magnetic ad-atom arrays that mimic magnetic chains, where spin-orbit coupling can enhance the AC phase [61].
Acknowledgments
@@
We acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC–2111–390814868, TRR 360 – 492547816, DFG grants No. KN1254/1-2, KN1254/2-1, and DFG Emmy Noether Programme—Project No. 504261060, the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 851161), as well as the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
Data and Informations availability
@@
Data, data analysis, and simulation codes are available upon reasonable request on Zenodo [62].
@@
Appendix A The Strong Anisotropy Limit – Chiral Magnon Bound States
@@
In the main text we emphasized how strongly bound eigenstates can be tuned through resonances with individual asymptotic chiral states. Having increased the bare interaction strength $\Delta$ to sufficiently large values these strongly localized modes will transition to form stable bound modes, separated in energy from the two-particle continuum, as shown in Fig. 5 (a). In total a number of $L_{y}$ bound state bands will experience this effect resembling the results for the single magnon spectrum. Each bound state band, however, is associated with twice the shift in momentum space. This is a result of the AC phase coupling to the physical spin charge of the excitation; the latter is double for the bound state compared to individual magnon excitations. In the spirit of the single magnon spectrum sufficient coupling between the wires, nonetheless, opens a bulk gap stabilizing two chiral bound state modes at the edges of the system, see Fig. 5 (b) - (d). Same as the single magnon edge modes these edge bound states are stable throughout the tested parameter regime. This is consistent with findings from Ref. [56, 57].
Appendix B Exact Diagonalization in Subspaces of Fixed Magnon Number
The conclusions drawn in our work are based on numerical results for single and two magnon spectra as well as characteristics of the corresponding wavefunctions. All presented results are obtained using exact diagonalization of the discussed Hamiltonian working either in a real space formulation or directly resolving the spectrum in momentum space. Our simulations thereby make use of magnon number conservation implied in Hamiltonian (3) of the main text due to $U(1)$-symmetry around the axis of spin quantization $\hat{z}$. Exploiting this symmetry allows us to diagonalize sectors of a given magnon number individually. In contrast to an exponential scaling of the total Hilbert space dimension ($\propto 2^{N}$) with the size of the system $N=L_{x}\times L_{y}$ these subsectors scale algebraically with the number of magnons, i.e. linear $\propto N$ for a single respectively quadratic $\propto N^{2}$ for two magnons. This beneficial scaling allows us to reach the system sizes required to study scattering between two magnons. For the extraction of scattering phases presented in Appendix C.1 it is, moreover, required to obtain two-magnon eigenstates for a given total momentum $K$. For this, we assume periodic boundary conditions in the system along the wire direction $\hat{x}$ and derive the momentum space representation of $\mathcal{H}$. The latter is diagonal in the quantum number $K$ and can be diagonalized for each $K$ independently.
Appendix C Characterization of the Many-Body Theory of Edge Modes
C.1 Scattering Phases from Two Magnon Spectra
@@
Our investigations on the edge theory in the main text are primarily based on the momentum dependent analysis of scattering phases. For this, we rely on a protocol consisting of the following points to extract scattering information from the numerically obtained spectra, also summarized in Fig. 5 (e).
@@
    Determine the momentum-dependent energy window $E_{1}(K),E_{2}(K)]$ possible for pairs of chiral magnons from the single magnon spectrum.
@@
    Filter the eigenspectrum for the eigenenergies within $E_{1}(K),E_{2}(K)]$.
@@
    From the fitted parameters $q,A,B$ compute the scattering matrix $S(q)=\frac{B}{A}$ as a function of relative momentum $q$.
@@
$S_{\mathrm{XXZ}}(K,q)=e^{i\phi_{\mathrm{XXZ}}(K,q)}$
$\phi_{\mathrm{XXZ}}(K,q)=\pi-i\log\Bigg[\dfrac{1-4\Delta e^{i(\frac{K}{2}+q)}+e^{iK}}{1-4\Delta e^{i(\frac{K}{2}-q)}+e^{iK}}\Bigg].$ (9)
A comparison between the XXZ predicted (bottom row) and the numerically extracted (top row) scattering phases $\phi$ is shown in Fig. 5 (f) - (h) for anisotropies $\Delta\in[0.0,0.4,1.0]$. We collect scattering phases for values of $K\in[-\dfrac{\pi}{4},\dfrac{\pi}{4}]$ and show the results as a function of the localization length $\ell$ of the eigenstates. To further discriminate the individual data points we highlight their relative momenta via color coding. The strongest deviations appear at small $\Delta$ for less localized states. This fits the semiclassical expectation of two more delocalized edge states having a reduced chance of scattering during their orbital motion at the boundary. In reality this enhanced leaking into the bulk, however, abets hybridization with bulk modes and the existence of resonances.
Figure 5: Chiral Bound States and Scattering Phase Analysis. a) Bound state bands separated in energy from the two magnon continuum for anisotropy $\Delta=1.25$ and uncoupled wires. b) - d) Increasing hopping strength between the wires $J_{y}\in\{0.25,0.50,0.75\}$ opens a bulk gap and stabilizes modes of chiral bound states. e) Analysis of the scattering phase consists of (1.) - (3.) selecting eigenstates within the chiral energy window for a given momentum sector $K$, (4.) fitting the states against a plane wave ansatz and (5.) compute the scattering matrix $S$. f) - h) Comparison of extracted (top row) and XXZ predicted (bottom row) scattering phases for different $K$ momenta and anisotropies $\Delta\in\{0.0,0.4,1.0\}$. The associated relative momenta are highlighted by color coding. Deviations are evident at low momenta ($\Delta=0.0$) and especially for less localized states.
In the main text we also extract the effective interaction for the scattering phase data. For this, we determine the slope of $\phi(q)$ in the linear regime around relative momenta $q=0$. We correct the results by additionally weighting them with the effective magnon mass $m^{*}$ obtained from the curvature of the single magnon dispersion, i.e. $m^{*}(K)=\dfrac{1}{\partial_{K}^{2}\varepsilon(K)}$. The result quantifies the one-dimensional scattering problem yielding the inverse product of effective magnon mass and the scattering length $a$ for the edge theory, which is described in the next section. To quantify deviations from the bare XXZ theory we extract the same quantity from the analytical XXZ scattering prediction (C.1)
$c(K)=\Bigg[-\frac{1}{m^{*}a}\Bigg](K)=\dfrac{4\Delta\cos\big(K\big)}{2\Delta-\cos\bigg(\dfrac{K}{2}\bigg)}.$ (10)